# Particle transport and deposition in wall-sheared thermal turbulence

### v1

@@ 3.2 Particle deposition on the wall
To quantify the particle deposition behavior, we first examine the p.d.f.s of their deposition locations along the bottom wall, as shown in figure 11. In our simulations, the particles adhere to the wall upon contact, and we then mark them as deposited and record the contact position as deposition locations. For particles with small $St$ (see figure 11a), the deposition locations are uniformly distributed along the bottom wall, aligning with the assumption of a spatially uniform particle concentration. For medium $St$ (see figure 11b), there is pronounced aggregation behavior for deposition positions along the bottom wall. With the increase of $Re_{w}$, the degree of aggregation weakens, possibly because wall shear causes the boundary layers to move horizontally, thereby dispersing the particles to deposit more evenly along the bottom wall. We also observed a correlation between the regions of particle accumulation on the bottom wall and the rising of hot plumes (compared to figure 2), as revealed in our previous findings (Xu et al., 2020) and independent research by Patočka et al. (2020, 2022). The accumulation of heavy particles beneath upwellings seems somewhat counterintuitive, as one might expect these particles to accumulate primarily beneath significant downwellings. For large $St$ (see figure 11c), the uniformity of the deposition pattern is restored, mainly attributed to the reduced influence of the carrier flow on these particles, resulting in an initial uniform distribution less affected by the carrier flow, unlike their low $St$ counterparts.
Figure 11: The p.d.f.s of particle deposition locations along the bottom wall with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a) ${d_{p}}=5$ $\mu$m, (b) ${d_{p}}=20$ $\mu$m, and (c) ${d_{p}}=80$ $\mu$m.
@@
Figure 12: Settling curves from 2-D simulations with $Ra=10^{8}$ at (a-c) $Re_{w}=0$, (d-f) $Re_{w}=2000$, and (g-i) $Re_{w}=6000$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m, and (c,f,i) ${d_{p}}=80$ $\mu$m. The insets in (a,d,g) magnify differences between our new model and the exponential law, and insets in (c,f,i) magnify differences between our new model and the linear law.
Figure 13: Settling curves from 3-D simulations with $Ra=10^{7}$ at (a-c) $Re_{w}=0$, (d-f) $Re_{w}=2000$, and (g-i) $Re_{w}=4000$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5$ $\mu$m, (b,e,h) ${d_{p}}=20$ $\mu$m, and (c,f,i) ${d_{p}}=80$ $\mu$m. The insets in (a,d,g) magnify differences between our new model and the exponential law, and insets in (c,f,i) magnify differences between our new model and the linear law.
Despite existing simple models to predict particle deposition rates for small and large $St$, challenges arise for medium $St$ due to the complex particle-turbulence interactions. Examining the settling curves obtained from simulation, we develop the following model that comprises a linear settling stage succeeded by a non-linear settling stage. We denote $t_{s}$ as the transient time from the linear to the non-linear stage. At the early settling stage, particles initially released in the lower domain and in the downwelling areas tend to settle. To illustrate this, we conducted a posterior examination and marked the initial positions of particles that settle during the linear stage, as shown in figures 14 (a-c). The particle drift equation (Patočka et al., 2022) gives $\mathbf{u}_{p}=\mathbf{u}_{f}+\mathbf{v}_{t}$ in the limit of $St\ll 1$, where $\mathbf{v}_{t}$ is the particle Stokes velocity. This equation indicates that particles subjected to updrafts settle at a speed of ${v_{t}}-{v_{f}}$, while those in downdrafts settle at ${v_{t}}+{v_{f}}$. Here, ${v_{f}}$ denotes the vertical flow speed at the particle’s position. At this stage, the settling particles decrease height almost monotonously, without recirculating in the convection cell, as shown in figures 15 (a-c). Assuming the same proportion of settling particles affected by the updraft in the lower domain and downdraft in the downwelling areas, then the numbers of $N_{\text{up}}$ and $N_{\text{down}}$ during $t\leq{t_{s}}$ are
@@
thus, at this settling stage, the deposition ratio ( i.e., the fraction of particles that have deposited on the wall relative to the initial number of particles laden in the flow system) is
@@
Figure 14: A posterior examination of the initial positions (black dots) for particles that have settled (a-c) during the linear stage and (d-f) during the non-linear stage, alongside the corresponding instantaneous temperature fields (contours) at (a,d) $Re_{w}=0$, (b,e) $Re_{w}=2000$, and (c,f) $Re_{w}=6000$, with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3 and diameter is fixed as ${d_{p}}=20$ $\mu$m. Note this figure does not represent a ”real” instantaneous of particle distribution within the convection cell.
Figure 15: Example particle trajectories for those settled (a-c) during the linear stage and (d-f) during the nonlinear stage at (a,d) $Re_{w}=0$, (b,e) $Re_{w}=2000$, and (c,f) $Re_{w}=6000$, with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3 and diameter is fixed as ${d_{p}}=20$ $\mu$m. The trajectories are colored according to the normalized suspension time of the particles. The blue circle marks the initial position and the red star marks the deposition position.
At the later non-linear settling stage, particles initially released in the roll center and upwelling ares tend to settle, as illustrated in figures 14 (d-f). The particles circulate within the large-scale roll and may escape the circulation due to small-scale irregularities produced by the emergence of new plumes (Patočka et al., 2020). With the increase of wall shear, the lateral motion increases, making the particles more likely to escape the circulation, as shown in figures 15 (d-f). Inspired by the stochastic model proposed by Denzel et al. (2023), we denote the time duration for a particle to complete the circulation as $t_{c}$ and the escape probability from the circulation as $\lambda_{f}$. The time decay rate of particle number $N$ in the convection cell is
@@
thus, at the non-linear settling stage, the deposition ratio is
@@
In short, we propose a new model that describes the settling process via an early linear stage and a later non-linear stage, written as
@@
Previously, in figures 12 and 13, we compared results from this newly developed model with 2-D and 3-D simulation results, respectively. Good agreement was observed for convective flows across various $Re_{w}$ and particles with a wide range of $St$. Although our original motivation was to propose a model to predict particle deposition rates for medium $St$, the simulation results indicate that our model is applicable for small and large $St$ as well. In figures 12 and 13, we also plot the predictions from the Patočka et al.’s model (denoted as $f$ distribution) (Patočka et al., 2020, 2022). The key idea of the $f$ distribution is to describe particle settling as $dN/dt=-N(t)/(fH/v_{t})$, leading to a settling curve determined by $N(t)/N_{0}=\exp[-v_{t}t/(Hf)]$. We observe a clear departure in the settling curve, mainly because their model adopts a unified exponential solution to describe the settling process, whereas DNS results indicate a linear regime in the settling curves (see also in Patočka et al., 2020, figure 3). As acknowledged by Patočka et al. (2020), the imperfect fit of the observed settling curves is due to the fact that $f$ should be a function of time rather than remaining constant. This time dependency of $f$ results in a mistfit between the $f$ distribution and the observed settling curves, with the largest error occurring for particles with a medium Stokes number. Nevertheless, Patočka et al.’s model is still useful for estimating the characteristic time of sedimentation as discussed later in this paper.
For our new model to be fully closed and predictive, we still need to provide accurate modeling of parameters $t_{s}$ and $C_{2}$. In figure 16, we plot $t_{s}$ and $C_{2}$ as a function of $St_{f}$ for various $Re_{w}$ in the 2-D cell with $Ra=10^{8}$. Here, we adopt $St_{f}$ rather than $St_{K}$ as the input parameter because $St_{f}$ is simpler to determine from the flow parameters, making it more convenient for modeling. In appendix B, we further examine the Rayleigh number dependence of such relationship with $Ra=10^{7}$ and $10^{9}$. We can see from figure 16(a), for medium $St_{f}$ in the range of $10^{-3}\leq St_{f}\leq 10^{-2}$, the parameter $t_{s}$ can be empirically estimated as
@@
It implies that $t_{s}$ is minorly influenced by $Re_{w}$ when the convective flow state remains unchanged (either in the LSC or zonal flow state). Practically, to empirically estimate $t_{s}$ using the above equation, we should first determine the flow states, then $t_{s}$ can be empirically expressed as a function of the Stokes number. This $t_{s}$ value is essentially an ensemble-averaged value for all the particles. As for the parameter $C_{2}$ (see figure 16b), although its original definition is just a combination of unknowns $t_{c}$ and $\lambda_{f}$ for the convenience of mathematical calculation, in the following, we will reveal the physical meaning of this parameter $C_{2}$.
@@
According to the particle mass conservation equation (Patočka et al., 2022; Martin & Nokes, 1989), the particle deposition rate $dN/dt$ is equal to the decreasing of particle flux $A_{0}v_{t}n_{\text{bnd}}$ at the lower boundary, namely, $dN/dt=-A_{0}v_{t}n_{\text{bnd}}$. Here, $A_{0}$ is the area (or length in 2-D case) of the bottom boundary, and $n_{\text{bnd}}$ is the mean particle number concentration near the lower boundary. Because the particle deposition rate is
@@
The above analysis indicates that the parameter $C_{2}$ describes the ratio of mean particle number concentration near the lower boundary over the volume-averaged particle concentration during the nonlinear stage. From figures 17(a-c), we observe good agreement on the $C_{2}$ values calculated via (31) with those determined by least square fitting of the settling curves. Then, in figures 17(d-f), we show the settling curves by determining unknown parameters via empirical correlation (29) for $t_{s}$ and (31) for $C_{2}$ to close the model. In addition to determining unknown parameters via fitting DNS data, here we highlight that using these relations provides an alternative solution to predict the settling curves. Recent work by Denzel et al. (2023) also relies on DNS data to determine the input statistics of their model. The difference is that in Denzel et al. (2023), the input parameters for their model are determined from the DNS data as a fit to the cubic splines of the discrete p.d.f.s. for those parameters, while the input parameters for our model are determined from the DNS data as a fit to the settling curves. More importantly, we also demonstrate predicting deposition ratio using empirical correlation (29) for $t_{s}$ and (31) for $C_{2}$ to close the model, which provides an alternative solution in addition to fitting DNS data to predict the settling curves.
Figure 17: (a-c) The parameter $C_{2}$ as a function of $St$, determined from data fitting of the settling curves and equation (31). (d-f) Settling curves obtained by determining unknown parameters via equation (29) for $t_{s}$ and (31) for $C_{2}$, shown for (a,d) $Re_{w}=0$, (b,e) $Re_{w}=2000$, and (c,f) $Re_{w}=6000$, with $Ra=10^{8}$. Particle density is fixed as $d_{p}$ = 3000 kg/m3 and diameter is fixed as $d_{p}$ = 20 $\mu$m.
@@
Figure 18: Particle mean residence time $\overline{{t_{\text{res}}}}$ normalized by the free-fall time $t_{f}$, with (a-c) $Ra=10^{7}$, (d-f) $Ra=10^{8}$, (g-i) $Ra=10^{9}$, at (a) $Re_{w}=0$, (b) $Re_{w}=1000$, (c) $Re_{w}=1500$, (d) $Re_{w}=0$, (e) $Re_{w}=2000$, (f) $Re_{w}=6000$, (g) $Re_{w}=0$, (h) $Re_{w}=9000$, and (i) $Re_{w}=12000$.
@@
For medium $St$, the particles circulate within the LSC, leading to a much longer average residence time compared to the average terminal time. With the increase of $Ra$, the LSC becomes more unstable (Zhu et al., 2018), and the locking of particles inside the LSC weakens progressively (Patočka et al., 2022). For example, at a low $Ra$ of $10^{7}$, the average residence time can be up to ten times longer than the average terminal time, due to a fraction of particles that are trapped inside the LSC, making the average value large; at a high $Ra$ of $10^{9}$, the average residence time is only slightly longer than the average terminal time. In addition, under the influence of wall shear, the LSC becomes unstable, leading to a decrease in the average residence time compared to the average terminal time. At $Ra=10^{8}$ and $Re_{w}=6000$ in the 2-D simulations, the updrafts in the zonal flow are much weaker due to the absence of uplifting force of plumes. Consequently, we observe faster settling of particles (see figure 19f).
Figure 19: Particle mean residence time $\overline{{t_{\text{res}}}}$ normalized by the mean terminal time $H/{v_{t}}$, with (a-c) $Ra=10^{7}$, (d-f) $Ra=10^{8}$, and (g-i) $Ra=10^{9}$, at (a) $Re_{w}=0$, (b) $Re_{w}=1000$, (c) $Re_{w}=1500$, (d) $Re_{w}=0$, (e) $Re_{w}=2000$, (f) $Re_{w}=6000$, (g) $Re_{w}=0$, (h) $Re_{w}=9000$, and (i) $Re_{w}=12000$.

### v2

@@ 3.2 Particle deposition on the wall
To quantify the particle deposition behaviour, we first examine the p.d.f.s of their deposition locations along the bottom wall, as shown in figure 11. In our simulations, the particles adhere to the wall upon contact, and we then mark them as deposited and record the contact position as the deposition location. For particles with small $St$ (see figure 11a), the deposition locations are uniformly distributed along the bottom wall, aligning with the assumption of a spatially uniform particle concentration. For medium $St$ (see figure 11b), there is pronounced aggregation behaviour for deposition positions along the bottom wall. With the increase of $Re_{w}$, the degree of aggregation weakens, possibly because wall shear causes the boundary layers to move horizontally, thereby dispersing the particles to deposit more evenly along the bottom wall. We also observed a correlation between the regions of particle accumulation on the bottom wall and the rising of hot plumes (compared with figure 2), as revealed in our previous findings (Xu et al., 2020) and independent research by Patočka et al. (2020, 2022). The accumulation of heavy particles beneath upwellings seems somewhat counterintuitive, as one might expect these particles to accumulate primarily beneath significant downwellings. For large $St$ (see figure 11c), the uniformity of the deposition pattern is restored, mainly attributed to the reduced influence of the carrier flow on these particles, resulting in an initial uniform distribution less affected by the carrier flow, unlike their low $St$ counterparts.
Figure 11: The p.d.f.s of particle deposition locations along the bottom wall with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a) ${d_{p}}=5\ \mu$m, (b) ${d_{p}}=20\ \mu$m and (c) ${d_{p}}=80\ \mu$m.
@@
Figure 12: Settling curves from 2-D simulations with $Ra=10^{8}$ at (a-c) $Re_{w}=0$, (d-f) $Re_{w}=2000$ and (g-i) $Re_{w}=6000$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5\ \mu$m, (b,e,h) ${d_{p}}=20\ \mu$m and (c,f,i) ${d_{p}}=80\ \mu$m. The insets in (a,d,g) magnify differences between our new model and the exponential law, and insets in (c,f,i) magnify differences between our new model and the linear law.
Figure 13: Settling curves from 3-D simulations with $Ra=10^{7}$ at (a-c) $Re_{w}=0$, (d-f) $Re_{w}=2000$ and (g-i) $Re_{w}=4000$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3, with diameters of (a,d,g) ${d_{p}}=5\ \mu$m, (b,e,h) ${d_{p}}=20\ \mu$m and (c,f,i) ${d_{p}}=80\ \mu$m. The insets in (a,d,g) magnify differences between our new model and the exponential law, and insets in (c,f,i) magnify differences between our new model and the linear law.
Despite existing simple models predicting particle deposition rates for small and large $St$, challenges arise for medium $St$ due to the complex particle-turbulence interactions. Examining the settling curves obtained from simulation, we develop the following model that comprises a linear settling stage succeeded by a nonlinear settling stage. We denote by $t_{s}$ the transient time from the linear to the nonlinear stage. At the early settling stage, particles initially released in the lower domain and in the downwelling areas tend to settle. To illustrate this, we conducted an a posterior examination and marked the initial positions of particles that settle during the linear stage, as shown in figures 14 (a-c). The particle drift equation (Patočka et al., 2022) gives $\mathbf{u}_{p}=\mathbf{u}_{f}+\mathbf{v}_{t}$ in the limit of $St\ll 1$, where $\mathbf{v}_{t}$ is the particle Stokes velocity. This equation indicates that particles subjected to updrafts settle at a speed of ${v_{t}}-{v_{f}}$, while those in downdrafts settle at ${v_{t}}+{v_{f}}$. Here, ${v_{f}}$ denotes the vertical flow speed at the particle’s position. At this stage, the settling particles decrease height almost monotonically, without recirculating in the convection cell, as shown in figures 15(a-c). Assuming the same proportion of settling particles affected by the updraft in the lower domain and downdraft in the downwelling areas, then the numbers of $N_{\text{up}}$ and $N_{\text{down}}$ during $t\leq{t_{s}}$ are
@@
thus, at this settling stage, the deposition ratio (i.e. the fraction of particles that have deposited on the wall relative to the initial number of particles laden in the flow system) is
@@
Figure 14: An a posterior examination of the initial positions (black dots) for particles that have settled (a-c) during the linear stage and (d-f) during the nonlinear stage, alongside the corresponding instantaneous temperature fields (contours) at (a,d) $Re_{w}=0$, (b,e) $Re_{w}=2000$ and (c,f) $Re_{w}=6000$, with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3 and diameter is fixed as ${d_{p}}=20$ $\mu$m. Note this figure does not represent a ’real’ instantaneous of particle distribution within the convection cell.
Figure 15: Example particle trajectories for those settled (a-c) during the linear stage and (d-f) during the nonlinear stage at (a,d) $Re_{w}=0$, (b,e) $Re_{w}=2000$ and (c,f) $Re_{w}=6000$, with $Ra=10^{8}$. Particle density is fixed as ${\rho_{p}}=3000$ kg/m3 and diameter is fixed as ${d_{p}}=20$ $\mu$m. The trajectories are coloured according to the normalized suspension time of the particles. The blue circle marks the initial position and the red star marks the deposition position.
At the later nonlinear settling stage, particles initially released in the roll centre and upwelling areas tend to settle, as illustrated in figures 14(d-f). The particles circulate within the large-scale roll and may escape the circulation due to small-scale irregularities produced by the emergence of new plumes (Patočka et al., 2020). With the increase of wall shear, the lateral motion increases, making the particles more likely to escape the circulation, as shown in figures 15(d-f). Inspired by the stochastic model proposed by Denzel et al. (2023), we denote the time duration for a particle to complete the circulation as $t_{c}$ and the escape probability from the circulation as $\lambda_{f}$. The time decay rate of particle number $N$ in the convection cell is
@@
thus, at the nonlinear settling stage, the deposition ratio is
@@
In short, we propose a new model that describes the settling process via an early linear stage and a later nonlinear stage, written as
@@
Previously, in figures 12 and 13, we compared results from this newly developed model with 2-D and 3-D simulation results, respectively. Good agreement was observed for convective flows across various $Re_{w}$ and particles with a wide range of $St$. Although our original motivation was to propose a model to predict particle deposition rates for medium $St$, the simulation results indicate that our model is applicable for small and large $St$ as well. In figures 12 and 13, we also plot the predictions from the Patočka et al. model (denoted as the $f$ distribution) (Patočka et al., 2020, 2022). The key idea of the $f$ distribution is to describe particle settling as $dN/dt=-N(t)/(fH/v_{t})$, leading to a settling curve determined by $N(t)/N_{0}=\exp[-v_{t}t/(Hf)]$. We observe a clear departure in the settling curve, mainly because their model adopts a unified exponential solution to describe the settling process, whereas DNS results indicate a linear regime in the settling curves (see also in Patočka et al., 2020, figure 3). As acknowledged by Patočka et al. (2020), the imperfect fit of the observed settling curves is due to the fact that $f$ should be a function of time rather than remaining constant. This time dependency of $f$ results in a misfit between the $f$ distribution and the observed settling curves, with the largest error occurring for particles with a medium Stokes number. Nevertheless, Patočka et al.’s model is still useful for estimating the characteristic time of sedimentation, as discussed later in this paper.
For our new model to be fully closed and predictive, we still need to provide accurate modelling of the parameters $t_{s}$ and $C_{2}$. In figure 16, we plot $t_{s}$ and $C_{2}$ as a function of $St_{f}$ for various $Re_{w}$ in the 2-D cell with $Ra=10^{8}$. Here, we adopt $St_{f}$ rather than $St_{K}$ as the input parameter because $St_{f}$ is simpler to determine from the flow parameters, making it more convenient for modelling. In Appendix B we further examine the Rayleigh number dependence of such a relationship with $Ra=10^{7}$ and $10^{9}$. We can see from figure 16(a), for medium $St_{f}$ in the range of $10^{-3}\leq St_{f}\leq 10^{-2}$, the parameter $t_{s}$ can be empirically estimated as
@@
This implies that $t_{s}$ is minorly influenced by $Re_{w}$ when the convective flow state remains unchanged (either in the LSC or zonal flow state). Practically, to empirically estimate $t_{s}$ using the above equation, we should first determine the flow states, then $t_{s}$ can be empirically expressed as a function of the Stokes number. This $t_{s}$ value is essentially an ensemble-averaged value for all the particles. As for the parameter $C_{2}$ (see figure 16b), although its original definition is just a combination of unknowns of $t_{c}$ and $\lambda_{f}$, for the convenience of mathematical calculation, in the following, we will reveal the physical meaning of this parameter $C_{2}$.
@@
According to the particle mass conservation equation (Patočka et al., 2022; Martin & Nokes, 1989), the particle deposition rate $dN/dt$ is equal to the decrease of particle flux $A_{0}v_{t}n_{\text{bnd}}$ at the lower boundary, namely, $dN/dt=-A_{0}v_{t}n_{\text{bnd}}$. Here, $A_{0}$ is the area (or length in the 2-D case) of the bottom boundary, and $n_{\text{bnd}}$ is the mean particle number concentration near the lower boundary. Because the particle deposition rate is
@@
The above analysis indicates that the parameter $C_{2}$ describes the ratio of mean particle number concentration near the lower boundary over the volume-averaged particle concentration during the nonlinear stage. From figure 17(a-c), we observe good agreement of the $C_{2}$ values calculated via (31) with those determined by least square fitting of the settling curves. Then, in figure 17(d-f), we show the settling curves by determining unknown parameters via empirical correlation (29) for $t_{s}$ and (31) for $C_{2}$ to close the model. In addition to determining unknown parameters via fitting DNS data, here, we highlight that using these relations provides an alternative solution to predicting the settling curves. Recent work by Denzel et al. (2023) also relies on DNS data to determine the input statistics of their model. The difference is that, in Denzel et al. (2023), the input parameters for their model are determined from the DNS data as a fit to the cubic splines of the discrete p.d.f.s. for those parameters, while the input parameters for our model are determined from the DNS data as a fit to the settling curves. More importantly, we also demonstrate prediction of the deposition ratio using empirical correlation (29) for $t_{s}$ and (31) for $C_{2}$ to close the model, which provides an alternative solution in addition to fitting DNS data to predict the settling curves.
Figure 17: (a-c) The parameter $C_{2}$ as a function of $St$, determined from data fitting of the settling curves and (31). (d-f) Settling curves obtained by determining unknown parameters via (29) for $t_{s}$ and (31) for $C_{2}$, shown for (a,d) $Re_{w}=0$, (b,e) $Re_{w}=2000$ and (c,f) $Re_{w}=6000$, with $Ra=10^{8}$. Particle density is fixed as $d_{p}$ = 3000 kg/m3 and diameter is fixed as $d_{p}$ = 20 $\mu$m.
@@
Figure 18: Particle mean residence time $\overline{{t_{\text{res}}}}$ normalized by the free-fall time $t_{f}$, with (a-c) $Ra=10^{7}$, (d-f) $Ra=10^{8}$, (g-i) $Ra=10^{9}$, at (a) $Re_{w}=0$, (b) $Re_{w}=1000$, (c) $Re_{w}=1500$, (d) $Re_{w}=0$, (e) $Re_{w}=2000$, (f) $Re_{w}=6000$, (g) $Re_{w}=0$, (h) $Re_{w}=9000$ and (i) $Re_{w}=12000$.
@@
For medium $St$, the particles circulate within the LSC, leading to a much longer average residence time compared with the average terminal time. With the increase of $Ra$, the LSC becomes more unstable (Zhu et al., 2018), and the locking of particles inside the LSC weakens progressively (Patočka et al., 2022). For example, at a low $Ra$ of $10^{7}$, the average residence time can be up to ten times longer than the average terminal time, due to a fraction of particles that are trapped inside the LSC, making the average value large; at a high $Ra$ of $10^{9}$, the average residence time is only slightly longer than the average terminal time. In addition, under the influence of wall shear, the LSC becomes unstable, leading to a decrease in the average residence time compared with the average terminal time. At $Ra=10^{8}$ and $Re_{w}=6000$ in the 2-D simulations, the updrafts in the zonal flow are much weaker due to the absence of the uplifting force of the plumes. Consequently, we observe faster settling of particles (see figure 19f).
Figure 19: Particle mean residence time $\overline{{t_{\text{res}}}}$ normalized by the mean terminal time $H/{v_{t}}$, with (a-c) $Ra=10^{7}$, (d-f) $Ra=10^{8}$ and (g-i) $Ra=10^{9}$, at (a) $Re_{w}=0$, (b) $Re_{w}=1000$, (c) $Re_{w}=1500$, (d) $Re_{w}=0$, (e) $Re_{w}=2000$, (f) $Re_{w}=6000$, (g) $Re_{w}=0$, (h) $Re_{w}=9000$ and (i) $Re_{w}=12000$.